# LifeGPT: Topology-Agnostic Generative Pretrained Transformer Model for Cellular Automata

### v1

@@ Abstract
The Game of Life (Life), a well known algorithm within the broader class of cellular automata (CA), exhibits complex emergent dynamics, with extreme sensitivity to initial conditions. Modeling and predicting such intricate behavior without explicit knowledge of the system’s underlying topology presents a significant challenge, motivating the development of algorithms that can generalize across various grid configurations and boundary conditions. We develop a decoder-only generative pretrained transformer model to solve this problem, showing that our model can simulate Life on a toroidal grid with no prior knowledge on the size of the grid, or its periodic boundary conditions (LifeGPT). LifeGPT is topology-agnostic with respect to its training data and our results show that a GPT model is capable of capturing the deterministic rules of a Turing-complete system with near-perfect accuracy, given sufficiently diverse training data. We also introduce the idea of an ‘autoregressive autoregressor’ to recursively implement Life using LifeGPT. Our results pave the path towards true universal computation within a large language model (LLM) framework, synthesizing of mathematical analysis with natural language processing, and probing AI systems for situational awareness about the evolution of such algorithms without ever having to compute them. Similar GPTs could potentially solve inverse problems in multicellular self-assembly by extracting CA-compatible rulesets from real-world biological systems to create new predictive models, which would have significant consequences for the fields of bioinspired materials, tissue engineering, and architected materials design.
@@ 1 Main
Cellular automata (CA) have long been a subject of profound interest within the fields of computer science and mathematics, owing to their intricate and emergent behaviors. CA algorithms are uniquely characterized by their combination of computational simplicity—evolving solely by local state-transition rules—and broad dynamical behavior, encompassing static, periodic, chaotic, and complex patterns, depending on the ruleset and initial condition (IC) being used. These properties render CA algorithms particularly valuable for simulating a wide array of natural phenomena, such as the propagation of forest fires[1], traffic flow dynamics[2], chemical reactions[3], and recrystallization[4]. The inherent behavioral unpredictability of CA (following human inspection of their rulesets) has hindered advancements in these subfields, confining CA to the realm of phenomenological modeling and subsequently preventing their evolution into mature, predictive tools for systems or phenomena where do not yet know a closed-form ruleset.
@@
One notable CA algorithm is the Game of Life (Life), as introduced by John H. Conway in 1970 [5]. Life (also sometimes referred to as ‘Conway’s Game of Life’) has since intrigued researchers from a range of disciplines, including mathematics, computer science and materials science. This is partly due to Life’s self-organizing yet often unpredictable dynamics[6] that emerge from a simple 2-state-8-neighbor transition scheme (Fig. 1). Despite Life’s fame, complex dynamics are present in numerous CA, and have led many researchers to support the conjecture that most CA are ‘computationally irreducible.’ This designation suggests that their evolution cannot be perfectly predicted by any process more computationally efficient than running the CA algorithms themselves[7]. Practically, this conjecture implies that the evolution of most CA algorithms from an arbitrary IC to an arbitrary time-step in the future cannot be described analytically. Thereby, any model that tries to simulate or predict the dynamics of a computationally irreducible CA is either implementing an algorithm capable of approximating CA behavior, or implementing the exact CA algorithm as a recursive operation for the number of timesteps desired. The former solution requires that the model diverge from the CA algorithm’s ground truth after some number of timesteps and/or for some specific initial conditions, and the latter solution necessitates that the model be fundamentally limited by its size to a maximum number of timesteps. Another question is whether, and how, we can begin to build artificially intelligent systems that can program new versions of CA to meet certain target behaviors, and whether such a system could develop a heightened situational ‘awareness’ to predict the evolution of an algorithm multiple timesteps into the future without explicitly computing it recursively. While this may seem impossible given the computational irreducibility conjecture, past research has suggested that Life exhibits self-organized-criticality (SOC)[6] and persistent scaling behavior [8]. Additionally, some CA have been shown to generate fractal-like patterns[9]. These phenomena suggest that Life (and potentially, many other CA algorithms) are at least statistically predictable – while perfectly accurate predictions of any CA system might be unachievable without running the exact CA algorithm, being able make global claims about CA systems is far more reasonable. The nexus question here is: Are there ‘deep trends’ that exist across, or transcend, scales in CA that might be elucidated by deep learning models? These and related questions led us to explore the use of AI algorithms in modeling CA.
@@
With increases in computing power and the growing popularity of neural networks as tools of analysis for complex systems (for which there are few practically useful analytical models), the computational irreducibly problem has lead many down the road of using neural networks to try to ‘predict’ the evolution of CA systems (often, Life is chosen as the example system), with varying degrees of success. The general anticipation across this field of research is that neural networks can either learn the rules of a given CA algorithm exactly, or can learn to abstract the system’s behavior well enough to find an accurate prediction of some far-future state based on an IC, though some work ventures beyond this paradigm.
Previous work focuses largely on either feed-forward convolutional neural networks (CNNs)[10] or convolutional encoder-decoder models[11]. Such approaches already encode knowledge of the significance of the spatial relationships between each cell in the CA training data, namely that CA utilize state-transition rules based on the neighboring states of a given cell (also known as the Moore neighborhood[12, 13]). CNNs embed the solution in the model directly because they already understand that CA utilize a 2D grid as a logical topology and hence even before the their weights are initialized, CNNs are more than half way to the solution. To the best of our knowledge, no research has investigated the possibility of utilizing a data topology-agnostic model for predicting the evolution of CA.
@@
In this paper we present an analysis of the transformer architecture’s capability to predict the outcomes of Life, played on a 32 $\times$ 32 toroidal grid, with remarkable accuracy. We accomplish this using a decoder-only transformer model, equipped with causally masked multi-headed self-attention, with forgetful causal masking (FCM) implemented during training, trained on data representing pairs of 2D grids encoding ICs and next-game-states (NGSs).
@@
ALife is a highly interdisciplinary field, focusing on understanding the essential attributes of life and life-like systems through computation, hardware, and biochemical means. ‘Soft’ ALife, which is the most relevant to our work, is a subset of ALife that is entirely simulated computationally, instead of being physically realized[35]. Many CA systems (notably, Life), have been constructed for the purpose of better understanding self-assembly, self-propagating patterns, and, more generally, complexity and chaos, all phenomenological attributes intrinsic to life as it is currently understood[36, 37, 38, 39].
Some scholars have already begun to conceptualize GPTs – more specifically, large language models (LLMs) – as having a ‘reciprocal relationship’ with ALife research. Nisioti et al. (2024)[40] identifies two key relationships between the two areas of research. The authors argue that LLMs may become tools for ALife research by allowing for the generation of open-ended environments, or as operators of evolutionary computation. Inversely, they also argue, ALife principles of such as collective intelligence, evolution, and organization may also be exhibited by LLM agents. The authors emphasize the potential for LLMs for control or understand ALife systems, which is the primary focus of our paper. While the scope of this paper is limited to a replication of Life’s rules using a transformer model, our work lays the groundwork for a paradigm in which GPTs are able to fine-tune CA systems’ emergent qualities through alterations of ICs and rulesets, enabling a form of ALife oversight and regulation, which the authors argue is a primary ethical consideration.
@@
Earlier work synthesizing the domains of CA and ML has largely focused on convolutional models paired with 2D CA. This is because 2D CA ‘time-slices’ are easily represented as images, which feed-forward CNNs are uniquely equipped to process. Springer and Kenyon (2021)[10] conducted a thorough empirical analysis on the ability of feed-forward CNNs to capture the allegedly computationally irreducible dynamics of Life. Since Life abides by rules pertaining to the states of 2D nearest-neighbors (the $r=1$ Moore neighborhood), which are effectively equivalent to a convolutional operations, the authors were able to calculate a hypothetical minimal feed-forward CNN (using ReLU activation) capable of capturing the rules of the Life and applying them over $n$ time-steps. They were thereby able to define a metric for network over-completeness, $m$, to empirically test the relationship between CNN size and Life learning effectiveness for various $n$ and $m$. One finding was that for $n\geq 3$, increasing $m$ up to 25 was insufficient for reaching a training convergence fraction of over 50%. Furthermore, the authors found that for networks which did converge, the number of epochs necessary increased substantially with $n$. The authors also note that converged, minimal networks were very sensitive to sign-flipping of initial weights, suggesting that for smaller networks, more luck in initializing suitable weights is needed to converge. The authors further reported that gradient descend was highly sensitive to the distribution parameters of the dataset, suggesting that some game examples were more useful that others for teaching the rules of Life to the model.
@@
Wulff and Hertz (1992) demonstrated the use of simple neural networks with short range connections ($\Sigma-\Pi$ networks) to learn the dynamics of 2D CA[41]. They concluded that even with a network sharing the same topology as the CA being studied, the dynamics of Life could not be learned effectively without weight-sharing between neurons. Their early approach was similar to many modern-day graph neural network architectures. Still, their approach encoded part of the solution to the problem in the architecture of the model. By enabling weight-sharing and employing short-range connections, the architecture of the network was itself a reflection of the grid-topology and nearest-neighbor rules engendering Life.
@@
Recent work has shown that CA systems can be trained to dynamically sustain desired patterns (such as small 2D images), by allowing a single CNN to simultaneously control the state-transition rules of all cells in the system, which is likewise made possible by allowing cells to take on continuous states (as opposed to discrete states). By training the neural network (through repeated runs of the CA growth process, following by subsequent loss calculation by comparison with the desired image, followed by backpropagation), the model’s parameters are tuned so that a single cell can, over time, multiply into a cellular collective that ‘grows’ into the desired image[42].
@@
The plan of the paper is as follows. First, we introduce the LifeGPT architecture and discuss the effectiveness of the model under a variety of training and inference strategies, draw conclusions regarding the manner in which the LifeGPT learns, and discuss zero/few-shot learning capabilities of LifeGPT. Next, we introduce the concept of the ‘autoregressive autoregressor’ (ARAR), which allows LifeGPT to recursively simulate Life’s dynamics over multiple time steps. Finally, we discuss the broader implications of LifeGPT for ALife, universal computation within an LLM framework, and solving inverse problems for the design of biomimetic and bioinspired materials. Looking ahead, we provide an argument for why enhacing LifeGPT with reinforcement learning (RL) techniques could improve inference accuracy, extend its abilities to a wide space of CA rulesets, and further enhance its utility in simulating and understanding complex, life-like systems.
@@ 2.1 Model Convergence During Training
Details on the implementation of the model, training and datasets used are provided in Materials and Methods (section 4). LifeGPT is architected as a decoder-only transformer model with causally-masked self attention – in addition to FCM during training – which utilized 12 transformer layers and 8 attention heads to capture complex patters in Life. The model’s maximum sequence length were configured to accommodate a data corresponding to ICs and NGSs. Rotary positional embedding (RPE) was used to maintain spatial awareness, while the Adam optimizer and cross-entropy loss (CEL) function guided the training process (see section 4.2 for a full list of hyperparameters). The best-performing version of LifeGPT was trained on training data with broad-entropy ICs, though using high-entropy ICs was also tested (see sections 4.3.5 and 2.3).
@@ 2.2 Accuracy Benchmarking and Sampling Temperature Effects
We observed temperature does have a substantial impact on accuracy (Fig. 2B). Even though the model was trained on deterministic data, it was prone to some minor errors, especially as sampling temperature is increased. Nevertheless, even with errors, models across all temperatures tested (0.0, 0.25, 0.50, 0.75, 1.0) still achieved at least 99.9% accuracy after about 20 epochs. Moreover, the variability in accuracy across temperatures was observed to decrease as the number of epochs of training were increased, suggesting that training for more epochs may further mitigate errors. For this task, a temperature of 0.0 showed the best empirically-measured accuracy scores across nearly all epochs. After epoch 16, the model rarely failed to achieve 100% accuracy, although there were couple exceptions where the model missed a one or two tokens. Indeed, even for the same model trained for 50 epochs, there was one sample in the testing set, ‘r-pentomino’ (see r-Pentomino Animation at Temperature 0 – 10 Timesteps), for which LifeGPT’s recursive NGS predictions failed to agree with the ground truth, due to a single incorrectly predicted cell early on.

### v2

@@ Abstract
Conway’s Game of Life (Life), a well known algorithm within the broader class of cellular automata (CA), exhibits complex emergent dynamics, with extreme sensitivity to initial conditions. Modeling and predicting such intricate behavior without explicit knowledge of the system’s underlying topology presents a significant challenge, motivating the development of algorithms that can generalize across various grid configurations and boundary conditions. We develop a decoder-only generative pretrained transformer (GPT) model to solve this problem, showing that our model can simulate Life on a toroidal grid with no prior knowledge on the size of the grid, or its periodic boundary conditions (LifeGPT). LifeGPT is topology-agnostic with respect to its training data and our results show that a GPT model is capable of capturing the deterministic rules of a Turing-complete system with near-perfect accuracy, given sufficiently diverse training data. We also introduce the idea of an ‘autoregressive autoregressor’ to recursively implement Life using LifeGPT. Our results pave the path towards true universal computation within a large language model framework, synthesizing of mathematical analysis with natural language processing, and probing AI systems for situational awareness about the evolution of such algorithms without ever having to compute them. Similar GPTs could potentially solve inverse problems in multicellular self-assembly by extracting CA-compatible rulesets from real-world biological systems to create new predictive models, which would have significant consequences for the fields of bioinspired materials, tissue engineering, and architected materials design.
@@ 1 Main
Cellular automata (CA) have long been a subject of profound interest within the fields of computer science and mathematics, owing to their intricate and emergent behaviors. CA algorithms are uniquely characterized by their combination of computational simplicity—evolving solely by local state-transition rules—and broad dynamical behavior, encompassing static, periodic, chaotic, and complex patterns, depending on the ruleset and initial condition (IC) being used. These properties render CA algorithms particularly valuable for simulating a wide array of natural phenomena, such as the propagation of forest fires [1], traffic flow dynamics [2], chemical reactions [3], and recrystallization [4]. The inherent behavioral unpredictability of CA (following human inspection of their rulesets) [5] has hindered advancements in these subfields, confining CA to the realm of phenomenological modeling and subsequently preventing their evolution into mature, predictive tools for systems or phenomena where do not yet know a closed-form ruleset.
@@
One notable CA algorithm is Conway’s Game of Life (Life), as introduced by John H. Conway in 1970 [6]. Life (also referred to as simply ‘the Game of Life’) has since intrigued researchers from a range of disciplines, including mathematics, computer science and materials science. This is partly due to Life’s self-organizing yet often unpredictable dynamics [7] that emerge from a simple 2-state-8-neighbor transition scheme (Fig. 1). Despite Life’s fame, complex dynamics are present in numerous CA, and have led many researchers to support the conjecture that most CA are ‘computationally irreducible.’ This designation suggests that their evolution cannot be perfectly predicted by any process more computationally efficient than running the CA algorithms themselves [5]. Practically, this conjecture implies that the evolution of most CA algorithms from an arbitrary IC to an arbitrary time-step in the future cannot be described analytically. As Wolfram (2024) [8] writes, “… we can’t expect to systematically ‘jump ahead” and predict […] the system…” Thereby, any model that tries to simulate or predict the dynamics of a computationally irreducible CA is either implementing an algorithm capable of approximating CA behavior, or implementing the exact CA algorithm as a recursive operation for the number of timesteps desired. The former solution requires that the model diverge from the CA algorithm’s ground truth after some number of timesteps and/or for some specific initial conditions, and the latter solution necessitates that the model be fundamentally limited by its size to a maximum number of timesteps. Another question is whether, and how, we can begin to build artificially intelligent systems that can program new versions of CA to meet certain target behaviors, and whether such a system could develop a heightened situational ‘awareness’ to predict the evolution of an algorithm multiple timesteps into the future without explicitly computing it recursively. While this may seem impossible given the computational irreducibility conjecture, past research has suggested that Life exhibits self-organized-criticality (SOC) [7] and persistent scaling behavior [9]. Additionally, some CA have been shown to generate fractal-like patterns [10]. These phenomena suggest that Life (and potentially, many other CA algorithms [11]) are at least statistically predictable – while perfectly accurate predictions of any CA system might be unachievable without running the exact CA algorithm, being able make global or ‘coarse-grained’ claims about CA systems is far more reasonable. In addition, recent work [12] demonstrates that exposing LLMs to complex elementary CA systems (class 4 [5]), as opposed to static (class 1), periodic (class 2), or chaotic (class 3) CA systems, leads to improved downstream performance on cognitive tasks, suggesting that a fundamental understanding of complexity is intrinsic to the common notion of ‘intelligence’. The nexus question here is: Are there ‘deep trends’ (which may also be conceptualized as ‘pockets of computational reducibility’ [8]) that exist across, or transcend, scales in CA that might be elucidated by deep learning models? These and related questions led us to explore the use of AI algorithms in modeling CA.
@@
With increases in computing power and the growing popularity of neural networks as tools of analysis for complex systems (for which there are few practically useful analytical models), the computational irreducibly problem has lead many down the road of using neural networks to try to ‘predict’ the evolution of CA systems (often, Life is chosen as the example system), with varying degrees of success. The general assumption across this field of research is that neural networks can either learn the rules of a given CA algorithm exactly, or can learn to abstract the system’s behavior well enough to find a more-or-less accurate prediction of some far-future state based on an IC, though some work ventures beyond this paradigm.
Previous work focuses largely on either feed-forward convolutional neural networks (CNNs) [13] or convolutional encoder-decoder models [14]. Such approaches already encode knowledge of the significance of the spatial relationships between each cell in the CA training data, namely that CA utilize state-transition rules based on the neighboring states of a given cell (also known as the Moore neighborhood [15, 16]). This likely introduces a form of inductive bias. CNNs embed the solution in the model directly because they already understand that CA utilize a 2D grid as a logical topology and hence even before the their weights are initialized, CNNs are more than half way to the solution. To the best of our knowledge, no research has investigated the possibility of utilizing a data topology-agnostic model for predicting the evolution of CA.
@@
In this paper we present an analysis of the transformer architecture’s capability to predict the outcomes of Life, played on a 32 $\times$ 32 toroidal grid, with remarkable accuracy. We accomplish this using a decoder-only transformer model, equipped with causally masked multi-headed self-attention, with forgetful causal masking (FCM) implemented during training, trained on data representing pairs of 2D grids encoding ICs and next-game-states (NGSs); we call our model LifeGPT.
@@
ALife is a highly interdisciplinary field that is focused on understanding the essential attributes of life and life-like systems through computation, hardware, and biochemical means. ‘Soft’ ALife, which is the most relevant to our work, is a subset of ALife that is entirely simulated computationally, instead of being physically realized [38]. Many CA systems (notably, Life), have been constructed for the purpose of better understanding self-assembly, self-propagating patterns, and, more generally, complexity and chaos, all phenomenological attributes intrinsic to life as it is currently understood [39, 40, 41, 42].
Some scholars have already begun to conceptualize GPTs – more specifically, large language models (LLMs) – as having a ‘reciprocal relationship’ with ALife research. Nisioti et al. (2024) [43] identify two key relationships between the two areas of research. The authors argue that LLMs may become tools for ALife research by allowing for the generation of open-ended environments, or as operators of evolutionary computation. Inversely, they also argued, ALife principles of such as collective intelligence, evolution, and organization may also be exhibited by LLM agents. The authors emphasize the potential for LLMs for control or understand ALife systems, which is the primary focus of our paper. While the scope of our paper is limited to a replication of Life’s game-state transition rules using a transformer model, our work lays the groundwork for a paradigm in which GPTs are able to fine-tune CA systems’ emergent qualities through alterations of ICs and rulesets, enabling a form of ALife oversight and regulation, which the authors argue is a primary ethical consideration.
Cellular Automata as Discrete Neural Network Models
In recent months, yet another reciprocal relationship between neural networks and ALife has been elucidated. Recent work by Wolfram [44] shows that discrete systems known as ‘spacetime-inhomogeneous cellular automata’ (also referred to as ‘rule arrays’) can be tuned through a discrete, Boolean-calculus-based backpropogation to create minimal models for perceptron-style neural-networks. In addition, it is argued that ordinary cellular automata can, similarly, discretely model recurrent neural networks. Through these examples, an argument is developed asserting that modern-day machine learning techniques are effective at learning simply because they can ‘mine’ programs from an extremely large space of complex behavior, arising from computationally irreducible systems governed by relatively few/simple rules. Thus, the author argues, it is not likely that AI will intrinsically favor finding ‘explainable’ or ‘understandable’ models for fitting training data, rather, AI will find something that just “happen[s] to work”. Furthermore, it is stated that in the context of AI development, this harnessing of computational irreducibly creates a fundamental trade-off between model interpretability/predictability and model capability. Models that are structured more obviously for human understanding will be more computationally reducible, and will therefore be forced to sample from a computational space with reduced complexity during weight-tuning. Moreover, it is suggested that models will intrinsically struggle to fit data generated by computationally irreducible programs, as finding programs that “do more or less the right thing” will not be sufficient. The work suggests that a deeper understanding of cellular automata will go hand-in-hand with the development of any kind of broadly applicable AI-science, as both necessitate the discovery and analysis of ‘pockets’ of computationally reducible behavior inside of complex systems.
@@
Earlier work synthesizing the domains of CA and ML has largely focused on convolutional models paired with 2D CA. This is because 2D CA ‘time-slices’ are easily represented as images, which feed-forward CNNs are especially equipped to process. Springer and Kenyon (2021) [13] conducted a thorough empirical analysis on the ability of feed-forward CNNs to capture the allegedly computationally irreducible dynamics of Life. Since Life abides by rules pertaining to the states of 2D nearest-neighbors (the $r=1$ Moore neighborhood), which are effectively equivalent to a convolutional operations, the authors were able to calculate a hypothetical minimal feed-forward CNN (using ReLU activation) capable of capturing the rules of the Life and applying them over $n$ time-steps. They were thereby able to define a metric for network over-completeness, $m$, to empirically test the relationship between CNN size and Life learning effectiveness for various $n$ and $m$. One finding was that for $n\geq 3$, increasing $m$ up to 25 was insufficient for reaching a training convergence fraction of over 50%. Furthermore, the authors found that for networks which did converge, the number of epochs necessary increased substantially with $n$. The authors also note that converged, minimal networks were highly sensitive to sign-flipping of initial weights, suggesting that for smaller networks, more luck in initializing suitable weights is needed to converge. The authors further reported that gradient descend was highly sensitive to the distribution parameters of the dataset, suggesting that some game examples were more useful that others for teaching the rules of Life to the model.
@@
Wulff and Hertz (1992) demonstrated the use of simple neural networks with short range connections ($\Sigma-\Pi$ networks) to learn the dynamics of 2D CA [45]. They concluded that even with a network sharing the same topology as the CA being studied, the dynamics of Life could not be learned effectively without weight-sharing between neurons. Their early approach was similar to many modern-day graph neural network architectures. Still, their approach encoded part of the solution to the problem in the architecture of the model, and in doing so introduced an inductive bias likely aided the learning process. By enabling weight-sharing and employing short-range connections, the architecture of the network was itself a reflection of the grid-topology and nearest-neighbor rules engendering Life.
@@
Recent work has shown that CA systems can be trained to dynamically sustain desired patterns (such as small 2D images), by allowing a single CNN to simultaneously control the state-transition rules of all cells in the system, which is likewise made possible by allowing cells to take on continuous states (as opposed to discrete states). By training the neural network (through repeated runs of the CA growth process, followed by subsequent loss calculation by comparison with the desired image and backpropagation), the model’s parameters are tuned so that a single cell can, over time, multiply into a cellular collective that ‘grows’ into the desired image [46].
@@
The plan of the paper is as follows. First, we introduce the LifeGPT architecture and discuss the effectiveness of the model under a variety of training and inference strategies, draw conclusions regarding the manner in which the LifeGPT learns, and discuss zero/few-shot learning capabilities of LifeGPT. Next, we introduce the concept of the ‘autoregressive autoregressor’ (ARAR), which allows LifeGPT to recursively simulate Life’s dynamics over multiple time steps. Finally, we discuss the broader implications of LifeGPT for ALife, universal computation within an LLM framework, and solving inverse problems for the design of biomimetic and bioinspired materials. Looking ahead, we provide an argument for why enhancing LifeGPT with reinforcement learning (RL) techniques could improve inference accuracy, extend its abilities to a wide space of CA rulesets, and further enhance its utility in simulating and understanding complex, life-like systems.
@@ 2.1 Model Convergence During Training
Details on the implementation of the model, training and datasets used are provided in Materials and Methods (section 4). LifeGPT is architected as a decoder-only transformer model with causally-masked self attention – in addition to FCM during training – which utilized 12 transformer layers and 8 attention heads to capture complex patterns in Life. The model’s maximum sequence length were configured to accommodate a data corresponding to ICs and NGSs. Rotary positional embedding (RPE) was used to maintain spatial awareness, while the Adam optimizer and cross-entropy loss (CEL) function guided the training process (see section 4.2 for a full list of hyperparameters). The best-performing version of LifeGPT was trained on training data with broad-entropy ICs, though using high-entropy ICs was also tested (see sections 4.3.5 and 2.3).
@@ 2.2 Accuracy Benchmarking and Sampling Temperature Effects
We observed temperature does have a substantial impact on accuracy (Fig. 2B). Even though the model was trained on deterministic data, it was prone to some minor errors, especially as sampling temperature is increased. Nevertheless, even with errors, models across all temperatures tested (0.0, 0.25, 0.50, 0.75, 1.0) still achieved at least 99.9% accuracy after about 20 epochs. Moreover, the variability in accuracy across temperatures was observed to decrease as the number of epochs of training were increased, suggesting that training for more epochs could have further mitigated errors. For this task, a temperature of 0.0 showed the best empirically-measured accuracy scores across nearly all epochs. After epoch 16, the model rarely failed to achieve 100% accuracy, although there were couple exceptions where the model missed one or two tokens. Indeed, even for the same model trained for 50 epochs, there was one sample in the testing set, ‘r-pentomino’ (see r-Pentomino Animation at Temperature 0 – 10 Timesteps), for which LifeGPT’s recursive NGS predictions failed to agree with the ground truth, due to a single incorrectly predicted cell early on.